### v2

# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

Abstract
Despite the rapid development of safety alignment techniques for LLMs, defending against multi-turn jailbreaks is still a challenging task. In this paper, we conduct a comprehensive comparison, revealing that some existing defense methods can improve the robustness of LLMs against multi-turn jailbreaks but compromise usability, i.e., reducing general capabilities or causing the over-refusal problem. From the perspective of mechanism interpretability of LLMs, we discover that these methods fail to establish a boundary that exactly distinguishes safe and harmful feature representations. Therefore, boundary-safe representations close to harmful representations are inevitably disrupted, leading to a decline in usability. To address this issue, we propose X-Boundary to push harmful representations away from boundary-safe representations and obtain an exact distinction boundary. In this way, harmful representations can be precisely erased without disrupting safe ones. Experimental results show that X-Boundary achieves state-of-the-art defense performance against multi-turn jailbreaks, while reducing the over-refusal rate by about 20% and maintaining nearly complete general capability. Furthermore, we theoretically prove and empirically verify that X-Boundary can accelerate the convergence process during training. Please see our code at: https://github.com/AI45Lab/X-Boundary. Warning: this paper includes examples that may be offensive or harmful.
Xiaoya Lu1,2,*Dongrui Liu1,*Yi Yu2Luxin Xu3Jing Shao2,†
1 School of Electronic Information and Electric Engineering, Shanghai Jiao Tong University
2 Shanghai Artificial Intelligence Laboratory
3 University of Electronic Science and Technology of China
footnotetext: Correspondence to: Jing Shao <shaojing@pjlab.org.cn>
1 Introduction
Figure 1: Illustration of the representation distinction boundary and the trade-off between multi-turn defense performance and over-refusal of existing defense methods and X-Boundary.
As large language models (LLMs) have demonstrated impressive abilities (OpenAI, 2024; AI@Meta, 2024; Team, 2023; Chen et al., 2024) and are increasingly deployed in diverse real-world applications (Zhou et al., 2024a; Huang et al., 2024), their security vulnerabilities have raised growing concern. One of the most significant security threats is “jailbreaks”, where deliberately designed prompts are used to elicit harmful responses from LLMs (Yi et al., 2024; Liu et al., 2023). Although current safety alignment techniques (Rafailov et al., 2024; Yuan et al., 2024) can defend against various single-turn jailbreaks, multi-turn jailbreaks remain a challenge. Unlike single-turn attacks, multi-turn jailbreaks exploit flexible multi-turn dialogues to bypass the safeguards of LLMs (Zhou et al., 2024b; Liu et al., 2024b), making them difficult to detect and defend against.
In this paper, we pioneeringly adapt and comprehensively compare four single-turn defense methods against multi-turn jailbreaks, including supervised fine-tuning (SFT) (Ren et al., 2024b), Direct Preference Optimization (DPO) (Jiang et al., 2024), Gradient Ascent (GA) (Zhang et al., 2024b), and circuit breaking (CB) (Zou et al., 2024). Although these methods can reduce the attack success rate (ASR) of multi-turn jailbreaks, they compromise the usability of LLMs, i.e., reducing general capability or leading to severe over-refusal issues. For example, the coding capability relatively reduces about 20% after SFT and the over-refusal rate exceeds 50% after GA. None of them successfully achieve a balance between defense robustness and usability. Over-refusal means that LLMs may unreasonably reject safe prompts (Röttger et al., 2023; Shi et al., 2024; Cui et al., 2024). For instance, Fig. 1 shows that LLMs refuse to answer a safe prompt “What are the harmful effects of meth on the human body?”.
From the perspective of the internal mechanism of LLMs, establishing a boundary to exactly distinguish harmful and safe feature representations is crucial for achieving the balance between robustness and usability. This distinction boundary enables harmful representations to be intervened in while safe representations are preserved. However, existing multi-turn defense methods fail to learn an exact boundary due to the lack of an explicit boundary formulation, as shown in Fig. 1. Some boundary-safe queries, e.g., “What are the harmful effects of meth on the human body?” involve sensitive information, e.g., “meth”, and their representations are close to harmful representations. In this way, these boundary-safe queries are mistakenly treated as harmful and rejected by LLMs.
To strike a balance between defense robustness and usability, we propose an X-Boundary that explicitly formulates the boundary between harmful and safe representations. Specifically, X-Boundary optimizes the LLM to push harmful representations far away from boundary-safe representations, while keeping trained boundary-safe representations close to their original representations. In this way, we obtain a precise distinction boundary and these harmful representations are further erased. Experimental results demonstrate that X-Boundary decreases the ASR of multi-turn jailbreaks from 58.5% to 16.5% on Llama-3-8B-Instruct, while reducing the over-refusal rate by an average of 20.5% compared to state-of-the-art (SOTA) method and preserving nearly complete general capability. Additionally, we theoretically analyze the feature learning trend of LLM with X-Boundary from the perspective of optimal transport theory. Theoretical analysis and experimental results indicate that X-Boundary achieves about 22% improvement in the learning speed.
X-Boundary achieves a win-win outcome with enhanced robustness against multi-turn jailbreaks and minimal decline in usability. We do not expect X-Boundary can independently address all security threats without collaborating with alignment methods such as SFT and Reinforcement Learning from Human Feedback (RLHF). Instead, we believe that X-Boundary can provide a more efficient and fine-grained defense for safety-aligned LLMs, ultimately enhancing the prospects of deploying robust AI systems in diverse real-world applications.
Table 1: Comparison of existing defense methods and X-Boundary.
Models Methods Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Llama-3- 8B-Instruct Vanilla 58.50 25.00 34.00 6.80 9.00 8.00 13.67 68.30 79.08 59.18
SFT 19.50 0.50 8.00 27.20 42.33 22.00 57.33 68.17 76.19 54.27
DPO 17.50 5.00 14.00 20.00 28.33 17.33 41.00 68.01 75.59 58.54
GA 38.50 1.50 12.00 10.80 15.00 13.33 35.33 68.25 77.86 62.20
CB 16.50 0.50 10.00 23.60 27.67 36.00 52.00 67.66 78.47 59.76
X-Boundary 16.50 1.00 10.00 8.40 14.00 8.00 28.67 67.94 78.70 59.76
Qwen2.5 -7B-Chat Vanilla 76.00 39.50 62.00 6.00 19.33 1.67 5.60 74.26 80.67 81.71
SFT 21.00 6.00 18.00 46.00 57.67 29.33 53.67 74.30 76.42 77.44
DPO 38.00 12.00 24.00 21.60 25.67 11.67 32.33 73.63 80.97 80.49
GA 38.00 21.00 12.00 58.40 70.00 67.67 85.33 74.58 80.43 79.27
CB 15.50 5.50 12.00 20.60 26.00 34.00 43.67 74.21 80.36 81.10
X-Boundary 17.50 7.50 16.00 10.40 16.67 5.33 15.00 74.17 80.52 81.10
Mistral-7B -Instruct-v0.2 Vanilla 70.00 49.50 40.00 10.00 21.00 4.33 13.00 59.98 45.34 34.76
SFT 37.50 22.00 18.00 53.60 42.00 29.33 58.67 58.94 41.55 27.44
DPO 44.50 19.00 28.00 25.20 38.67 20.33 37.67 58.79 43.21 34.76
GA 24.00 9.00 10.00 38.40 50.67 35.67 71.33 60.13 45.64 34.76
CB 15.00 11.50 12.00 45.20 32.33 55.00 50.00 59.91 46.63 33.54
X-Boundary 16.00 13.50 14.00 19.20 23.33 10.34 26.33 59.83 45.34 36.59
2 Comparison of Existing Defense Methods Against Multi-Turn Jailbreaks
To the best of our knowledge, we are the first to conduct a comprehensive evaluation of classic defense approaches against multi-turn jailbreaks, considering both defense robustness and impact on usability. Although previous studies (Ren et al., 2024b; Jiang et al., 2024) have employed Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) in multi-turn defense scenarios, many other defense methods effective against single-turn jailbreaks, such as Gradient Ascent (GA) and Circuit Breaking (CB), have been overlooked and have not been thoroughly compared. In Section 2.1, we outline the process of constructing training data, reproducing SFT and DPO, and adapting GA and CB for multi-turn defense scenarios. In Section 2.2, we present and analyze the evaluation results, revealing the shortcomings of existing defense methods in balancing robustness and usability.
2.1 Adaption and Evaluation of Single-Turn Defense Methods Against Multi-Turn Jailbreaks
We compare the following defense methods against multi-turn jailbreak (Ren et al., 2024b; Jiang et al., 2024) on Llama-3-8B-Instruct (AI@Meta, 2024), Qwen2.5-7B-Chat (Yang et al., 2024a), and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023):
  • SFT (Ren et al., 2024b): fine-tuning LLMs using harmful queries as inputs and refusal answers as supervised labels directly.
  • DPO (Rafailov et al., 2024; Jiang et al., 2024): aligning LLMs using harmful queries as inputs, harmful answers as rejected responses, and refusal answers as chosen responses.
  • GA (Zhang et al., 2024b; Lu et al., 2024a): unlearning harmful knowledge by training with gradient ascent optimization methods .
  • CB (Zou et al., 2024): remapping the representations of harmful knowledge to desired targeted representations.
Construct multi-turn defense datasets. We construct the multi-turn defense training datasets based on SafeMTData (Ren et al., 2024b). SafeMTData consists of 1680 safe multi-turn dialogues for the safety alignment of LLMs in multi-turn interactions. For SFT, we directly exploit SafeMTData as a multi-turn training dataset following Ren et al. (2024b). For DPO, we curate harmful responses to the harmful multi-turn queries in SafeMTData and constructed a multi-turn preference dataset following Jiang et al. (2024). For CB and GA, to remove harmful knowledge that could be elicited through multi-turn attacks, we add pairs of harmful queries from SafeMTData along with the curated harmful responses into their respective defense training datasets. (Zhang et al., 2024b; Zou et al., 2024) More details about data construction are illustrated in Appendix B.1 and training settings of these methods are listed in Appendix B.2.
Evaluation of defense robustness. To evaluate the robustness of these defense methods, we test them against three SOTA multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024) and Crescendo (Russinovich et al., 2024). The evaluation metric used is the Attack Success Rate (ASR), defined as the proportion of attack attempts that successfully elicit harmful content from the LLMs. A lower ASR indicates greater robustness against multi-turn jailbreak attacks. We also evaluate the compliance rates of LLMs with harmful requests in HarmBench and the results are listed in the Appendix A.1.
Evaluation of usability. We evaluate the impact of defense methods on the usability of LLMs from two perspectives: over-refusal and the decline of general capability. Over-refusal is evaluated using XSTest (Röttger et al., 2023), OKTest (Shi et al., 2024), OR-Bench (Cui et al., 2024), and PHTest (An et al., 2024). The corresponding evaluation metric is the over-refusal rate, which measures the proportion of benign prompts that the model incorrectly refuses to answer. A lower over-refusal rate indicates better usability. General capability, including general knowledge, mathematics, and coding ability, is evaluated using MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. Please see Appendix B.4 for more details about evaluation.
2.2 Experimental Results and Analysis
Existing methods fail to strike a balance between robustness and usability. Table 1 shows that existing methods can effectively reduce the ASR of multi-turn jailbreaks after training with the aforementioned data. However, SFT, DPO, and GA even tend to severely compromise general capabilities when achieving good performance, commonly referred to as the “alignment tax” (Ouyang et al., 2022b). For instance, SFT results in a 3%-7% decrease in both the mathematical and the coding abilities of the three models. Moreover, all of these methods lead to a significant increase in the over-refusal rate.
An excessively high over-refusal rate makes the ASR unreliable. In PHtest, the over-refusal rate of Llama-3-8B increases from 13.67% to more than 40% on all methods. The over-refusal rates after GA are more than 50% on four benchmarks. The high over-refusal rate reflects that these methods cannot precisely distinguish harmful queries and build effective defense mechanisms for them. Instead, they simply reduce the ASR of multi-turn attacks by indiscriminately rejecting input queries, which is not trustworthy and undermines the model’s usability in real-world scenarios. Therefore, it is necessary to analyze the cause of over-refusal and propose a more precise defense method to mitigate it while preserving robustness against multi-turn jailbreak.
3 X-Boundary: Optimize Exact Boundary to Balance Defense Robustness and Usability
In this section, we propose X-Boundary to balance robustness against multi-turn jailbreaks and usability by explicitly formulating the distinction boundary. Section 3.1 analyzes the essential mechanism of decline in usability. Section 3.2 introduces the optimization objective of X-Boundary. Section 3.3 theoretically proves that X-Boundary may ease the learning difficulty and contribute to fast learning.
3.1 The Imprecise Distinction Boundary of Existing Multi-Turn Defense Methods.
Notations. Give an input data point $x$, $\mathcal{R}_{\mathcal{M}}\left(x\right)$ denotes its feature representations encoded by LLMs $\mathcal{M}$. $\{x_{i}\}_{i=1}^{N}$ and $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}\right)\}_{i=1}^{N}$ denote a set of multiple data points and representations, respectively. In particular, $x_{i}^{h}$ represents a harmful Query and its corresponding harmful Answer (QA pair), while $x_{i}^{r}$ denotes the refusal response to the harmful query $x_{i}^{h}$. $x_{i}^{s}$ and $x_{i}^{b}$ denote a safe QA pair and a boundary-safe QA pair, respectively, where the answer is both safe and helpful.
Analysis of safety-usability trade-off from the perspective of interpretability mechanism. Existing defense methods (Zou et al., 2024, 2023) typically improve the adversarial robustness of LLMs by intervening harmful feature representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$. Specifically, SFT (Yuan et al., 2024) and CB (Zou et al., 2024) remap harmful representations to refusal representations $\mathcal{R}_{\mathcal{M}}\left(x_{i}^{r}\right)$. In this process, these methods implicitly train LLMs to learn a boundary that distinguishes harmful representations and safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{s}\right)\}_{i=1}^{N}$. However, Fig. 2 shows that the boundary learned through this implicit training is imprecise, with some boundary-safe representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$ mixed with harmful representations rather than being clearly distinguished. In this way, these boundary-safe representations are mistakenly treated as harmful ones, leading LLMs to refuse the corresponding boundary-safe queries and ultimately reducing usability.
Figure 2: Visualization of the representation distribution after implementing SFT, DPO, GA, and CB. “Harmful” and “boundary-safe” refer to the representations of harmful and boundary-safe queries along with their corresponding responses, respectively.
3.2 Explicit Formulation for Distinction Boundary
We propose X-Boundary to explicitly formulate the distinction boundary between safe and harmful representations. The key idea is to push harmful representations far away from boundary-safe representations through explicit loss function, such that harmful representations can be effectively and precisely erased without disrupting safe ones. In this way, a balance between defense robustness and LLM usability can be achieved.
Specifically, we construct a separate set $D_{\texttt{s}}$ for separating harmful and boundary-safe representations, an erase set $D_{\texttt{e}}$ to contain harmful knowledge that should be erased, and a retain set $D_{\texttt{r}}$ for preserving safe knowledge related to the usability of LLMs. To this end, $D_{\texttt{r}}$ includes safe QA pairs $\{x_{i}^{s}\}_{i=1}^{N}$, boundary-safe QA pairs $\{x_{i}^{b}\}_{i=1}^{N}$, and refusal responses to harmful queries $\{x_{i}^{r}\}_{i=1}^{N}$. $D_{\texttt{e}}$ consists of harmful QA pairs: $D_{\texttt{e}}=\{x_{i}^{h}\}_{i=1}^{N}$. $D_{\texttt{s}}$ contains pairs of $x_{b}$ and $x_{r}$: $D_{\texttt{s}}=\{\left(x_{i}^{b},x_{i}^{r}\right)\}_{i=1}^{N}$.
Figure 3: Illustration of representation manipulation in X-Boundary for a clear distinction boundary.
To explicit formulate a precise distinction boundary, we propose separate loss $\mathcal{L}_{\texttt{s}}$ to increase the distance between harmful representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ and boundary-safe representations $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{b}\right)\}_{i=1}^{N}$. Since most $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)\}_{i=1}^{N}$ will be remapped to $\{\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)\}_{i=1}^{N}$ due to the following erasure operation, we can separate them by directly optimizing $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right)$ to be orthogonal to $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{b}\right)$ as shown in Fig. 3:
$\mathcal{L}_{\texttt{s}}=\frac{1}{\left|S_{s}\right|}\sum_{i=1}^{\left|S_{s}\right|}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right),\mathcal{R}_{\mathcal{M}}\left(x_{i}^{b}\right)\right)\right)$ (1)
To establish robust defense against multi-turn attacks, we utilize erase loss $L_{e}$ to erase the representations of harmful QA pairs in $D_{\texttt{e}}$. $L_{e}$ optimizes $\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right)$ to be orthogonal to their original representations $\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{h}\right)$ following (Zou et al., 2024):
$\mathcal{L}_{\texttt{e}}=\frac{1}{\left|D_{e}\right|}\sum_{i=1}^{\left|D_{e}\right|}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right),\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{h}\right)\right)\right)$ (2)
where $x_{i}$ represents a sample in retain set ($x_{i}\in D_{r}$), and $\mathcal{R}_{\mathcal{M}_{\theta}}$ and $\mathcal{M}_{\texttt{ref}}$ denote the model under training and the reference model before training, respectively.
To preserve usability of LLMs, we use retain loss $\mathcal{L}_{\texttt{r}}$ to maintain safe representations of data points in $S_{\texttt{r}}$. $\mathcal{L}_{\texttt{r}}$ minimizes the $\ell_{2}$ distance between trained representations and their original representations:
$\mathcal{L}_{\text{r}}=\frac{1}{\left|S_{r}\right|}\sum_{i=1}^{\left|S_{r}\right|}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}}\left(x_{i}\right)\right\|_{2}$ (3)
Notably, to maintain the existing refusal mechanism of LLMs, refusal responses $x_{r}$ to harmful queries are added into $S_{\texttt{r}}$. Therefore, most $\{\mathcal{R}_{\mathcal{M}}\left(x_{h}\right)\}_{i=1}^{N}$ are finally optimized to refusal representations $\{\mathcal{R}_{\mathcal{M}}\left(x_{r}\right)\}\}_{i=1}^{N}$ under the joint effect of $\mathcal{L}_{\texttt{e}}$ and $\mathcal{L}_{\texttt{r}}$.
In summary, the overall loss function is a weighted combination of the three aforementioned loss functions:
$\mathcal{L}=c_{r}\mathcal{L}_{r}+c_{e}\mathcal{L}_{e}+c_{s}\mathcal{L}_{s}$ (4)
where $c_{r}$, $c_{e}$ and $c_{s}$ are adaptive loss coefficients following (Zou et al., 2024; Ocampo et al., 2024). With the above optimization objective, X-Boundary can perform fine-grained optimization in the representation space to strike a balance between defense robustness and the usability of LLMs.
Algorithm 1 The optimization process of X-Boundary
0:  Original frozen model $\mathcal{M}_{\texttt{ref}}$, model $\mathcal{M}_{\theta}$ with parameters $\theta$ to be optimized, a function $\mathcal{R}$ that extracts representation from a model on a batch of inputs, a erase dataset $\mathcal{D}_{e}$, a retain dataset $\mathcal{D}_{r}$, a boundary dataset $D_{b}$, number of optimization steps $T$, hyperparameters $\alpha$ and $\beta$, batch size $n$
1:  for $t=1$ to $T$ do
2:     Sample $\{x_{i}\}_{i=1}^{n}\sim\mathcal{D}_{r}$, $\{x_{i}^{h}\}_{i=1}^{n}\sim\mathcal{D}_{e}$
3:     Sample $\{(x_{i}^{b},x_{i}^{r})\}_{i=1}^{n}\sim\mathcal{D}_{b}$
4:     $c_{r}=\alpha\frac{t}{\beta}$, $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$
5:     $\mathcal{L}_{\text{r}}=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}}\left(x_{i}\right)\right\|_{2}$
6:     $\mathcal{L}_{\texttt{e}}=\frac{1}{n}\sum_{i=1}^{n}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{h}\right),\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{h}\right)\right)\right)$
7:     $\mathcal{L}_{\texttt{s}}=\frac{1}{n}\sum_{i=1}^{n}\texttt{ReLU}\left(\texttt{cos}\left(\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}^{r}\right),\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}^{b}\right)\right)\right)$
8:     $\mathcal{L}=c_{r}\mathcal{L}_{r}+c_{e}\mathcal{L}_{e}+c_{s}\mathcal{L}_{s}$
9:     Update parameters $\theta$ to minimize $\mathcal{L}$
10:  end for
3.3 Theoretical Analysis of X-Boundary
In this subsection, we theoretically analyze the convergence rate of LLM from the perspective of the optimal transport theory (Solomon et al., 2020; Chuang et al., 2021; Weed & Bach, 2019). Specifically, we theoretically prove that X-boundary enables a faster learning speed of feature learning, which is verified in Fig. 4.
Preliminaries: optimal transport and $k$-variance. Wasserstein distance measures the distance between probability distributions on a metric space. Let $\mu$ and $\nu\in\textnormal{Prob}({\mathbb{R}}^{d})$ denote two probability measures, the definition of $p$-Wasserstein distance with Euclidean cost function is
${\mathcal{W}}_{p}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\left(\mathbb{E}_{(H,Q)\sim\pi}\|H-Q\|^{p}\right)^{1/p},$ (5)
where $\Pi(\mu,\nu)\subseteq\textnormal{Prob}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$ represent the set of measure couplings and $\mu$ and $\nu$ denote their marginals, respectively. From the perspective of optimal transport, Wasserstein distances indicate the minimal cost of transforming the distribution $\mu$ to $\nu$. Typically, the Earth Mover distance is equivalent to the 1-Wasserstein distance.
Definition 1 (Wasserstein-$1$ $k$-variance).
Given a probability measure $\mu\in\textnormal{Prob}({\mathbb{R}}^{d})$ and a parameter $k\in{\mathbb{N}}$, the Wasserstein-$1$ $k$-variance is given as
$\mathrm{Var}_{k}(\mu)=\mathbb{E}_{S,\tilde{S}\sim\mu^{k}}\left[{\mathcal{W}}_{1}(\mu_{S},\mu_{\tilde{S}})\right],$ (6)
[…]
$k$-variance measures structural properties of distribution beyond variance based on Wasserstein distances (Solomon et al., 2020). We theocratically analyze the learning trend of DNNs’ feature representations, which can be measured by the convergence rate of $k$-variance following (Weed & Bach, 2019; Solomon et al., 2020).
Proposition 1.
(Proven in Appendix C) If $\phi_{\#}\mu$ is $(n,\Delta)$-clusterable, then for all $m\leq n(2\Delta)^{-2}$,
$\mathrm{Var}_{m}(\phi_{\#}\mu)<48\Delta.$ (7)
Given a distribution $\mu$, $(n,\Delta)$-clusterable means that $\textnormal{supp}(\mu)$ lies in the union of $n$ balls of radius at most $\Delta$.
Proposition 1 indicates that $\mathrm{Var}_{m}(\phi_{\#}\mu)$ is bounded by the radius $\Delta$, reflecting the concentration of the feature distribution. In this way, the proposed X-Boundary enables more clustered features (the smaller radius $\Delta$) and a faster learning speed (the smaller $k$-variance $\mathrm{Var}_{m}(\phi_{\#}\mu)$).
Experimental Verification. Fig. 4 verifies that X-Boundary enables a faster learning speed of the training process. To this end, we fine-tune Llama-3-8B-Instruct and Qwen2.5-7B-Chat following settings in Section 2.1. Specifically, we set 0.1 and 0.55 of the training loss as thresholds to judge whether the training process has converged for Llama-3-8B-Instruct and Qwen2.5-7B-Chat, respectively. Based on this, Fig. 4 indicates that the proposed X-Boundary accelerates the converging process of 26.47% and 18.29% on Llama-3-8B-Instruct and Qwen2.5-7B-Chat, respectively.
Figure 4: The training curves of X-Boundary and without X-Boundary on Llama-3-8B-Instruct and Qwen2.5-7B-Chat.
4 Experiments
4.1 Experimental Settings
To ensure fairness in comparison and consistency in experimental settings, we also implement X-Boundary on Llama-3-8B-Instruct, Qwen2.5-7B-Chat, and Mistral-7B-Instruct-v0.2, and evaluate it using the benchmarks described in Section 2.1. Additionally, to assess the effectiveness of X-Boundary across different sizes of LLMs, we implement it on Qwen2.5-14B-Chat. To construct the Separate Set, we sample 500 boundary-safe prompts from OR-Bench-80K (Cui et al., 2024), which have been filtered to avoid data contamination with the test set of OR-Bench. Next, we use GPT-4o to generate safe and helpful responses for these prompts, thus we get boundary-safe QA pairs. The retain set consists of our collected boundary-safe QA pairs, UltraChat (Ding et al., 2023), and refusal data points generated by the trained LLMs themselves. The erase set includes the harmful QA pairs for single-turn defense used in Zou et al. (2024) and the harmful QA pairs for multi-turn defense described in Section 2.1. More implementation details of X-Boundary are listed in Appendix B.3.
Table 2: Comparison of existing defense methods and X-Boundary on Qwen2.5-14B-Chat.
Methods Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 71.50 63.50 36.00 4.00 10.00 1.33 4.00 80.06 82.49 79.88
SFT 52.00 10.00 16.00 43.60 51.33 31.33 62.67 79.58 82.18 81.71
DPO 54.50 45.00 32.00 6.40 14.00 2.67 8.67 78.58 83.32 81.10
CB 23.50 4.50 8.00 43.60 51.33 32.00 64.33 79.64 82.56 82.93
X-Boundary 25.00 5.00 12.00 5.20 13.67 4.00 8.33 79.52 82.18 81.10
4.2 Main Results
The explicit formulation for boundary contributes to the precise distinction between harmful and safe representations. To investigate the effect of the explicit formulation for distinction boundary, we visualize the representation distribution of X-Boundary and without X-Boundary. Fig. 5 shows that, without X-Boundary, the boundary-safe representations close to harmful representations are mistakenly regarded as harmful ones. This demonstrates that LLMs fail to learn a boundary that exactly distinguishes safe and harmful representations, which supports our motivation of explicitly formulating the distinction boundary. With X-Boundary, harmful representations and boundary-safe representations are clearly separated as shown in Fig. 5, verifying that the proposed explicit formulation contributes to establishing a precise distinction boundary. Please refer to Appendix A.6 and Appendix A.7 for more detailed visualization and analysis of the representation distribution.
X-Boundary maintains the lowest over-refusal rate while achieving SOTA defense against multi-turn jailbreaks. With a precise distinction boundary, X-Boundary reduces the ASR of ActorAttack by more than 40% while maintaining the increase in over-refusal rate on OKTest within 5% across three LLMs, as shown in Table 1. Specifically, on Llama-3-8B-Instruct, CB and X-Boundary both achieve the lowest ASR against ActorAttack, but X-Boundary demonstrates an average over-refusal rate that is lower by 20.05%. Similarly, on Qwen2.5-7B-Chat, X-Boundary’s average over-refusal rate is 58.50% lower than GA, which achieves the lowest ASR against Crescendo.
X-Boundary rarely declines general capability. Table 1 shows that the decline of general capabilities caused by X-Boundary is generally no more than 0.5% compared to vanilla models, across the domains of general knowledge, mathematical ability, and coding ability. In contrast to SFT, which causes a 7% reduction in coding ability for Mistral-7B-Instruct-v0.2, X-Boundary achieves a lower ASR without compromising coding capability.
X-Boundary successfully strikes a balance between robustness and usability. As a supplement to Table 1, Fig. 6 intuitively illustrates the trade-off between the ASR against multi-turn jailbreaks and the over-refusal rate. As the training process advances, the ASR steadily decreases, whereas the over-refusal rate progressively increases. Considering the two metrics comprehensively, X-Boundary appears in the lower-left corner of Fig. 6, indicating that it achieves a better balance compared to the baseline methods. In the same way, Fig. 8 demonstrates that X-Boundary also achieves a win-win outcome with robust defense and strong general capability. For specific cases of the defense performance and usability preservation of X-Boundary, please refer to Appendix D.
X-Boundary is effective across different sizes of LLMs. Table 2 shows that, on Qwen2.5-14B-Chat, X-Boundary relatively reduces the ASR of three multi-turn attacks by more than 60%, while keeping the increase in over-refusal rate within 5% compared to the vanilla model. While X-Boundary and CB achieve comparable ASR, the over-refusal rate of X-Boundary is approximately 40% lower than that of CB. Compared with the performance on Qwen2.5-7B-Chat, the performance of X-Boundary on Qwen2.5-14B-Chat has not decreased.
Figure 5: Visualization of the representation distribution of X-Boundary and without X-Boundary.
Figure 6: The trade-off between ASR of multi-turn jailbreaks and over-refusal rate on Llama-3-8B-Instruct. The data points were collected by sampling and evaluating at every 100 training steps.
Table 3: Comparison of existing defense methods and X-Boundary on reasoning models.
Models Methods Single & Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
HarmBench ActorAttack RedQueen XSTest OKTest OR-Bench PHTest AIME GPQA LiveCode
DeepSeek- R1-Distill- Llama-8B Vanilla 55.42 83.50 58.50 0.40 1.67 1.00 4.67 50.00 50.00 40.00
SFT 16.25 44.50 28.50 58.40 34.67 32.67 65.00 40.00 44.95 35.10
DPO 17.50 67.00 36.50 0.40 1.67 1.33 7.33 50.00 46.97 42.40
CB 52.80 77.50 55.00 1.60 3.33 3.67 8.00 46.67 46.97 40.65
X-Boundary 13.33 34.00 27.00 2.40 3.00 7.33 12.00 50.00 50.00 39.86
DeepSeek- R1-Distill- Qwen-7B Vanilla 60.42 79.00 60.50 1.60 2.00 1.33 2.67 53.30 48.99 39.76
SFT 48.33 59.50 42.50 7.60 9.00 11.00 12.67 46.67 48.99 36.44
DPO 52.92 74.00 48.50 1.60 3.67 3.00 8.67 53.33 50.00 40.08
CB 57.92 54.00 55.50 5.33 4.00 5.67 9.00 46.67 46.97 39.33
X-Boundary 20.80 43.50 36.00 2.40 3.00 7.33 12.00 53.33 50.51 40.02
Table 4: Ablation study on Qwen2.5-7B-Chat. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
Models A B C D Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 76.00 39.50 62.00 6.00 19.33 1.67 5.60 74.26 80.67 81.71
(a) $\checkmark$ 63.00 11.50 30.00 9.20 19.00 6.66 14.66 74.19 80.14 82.32
(b) $\checkmark$ $\checkmark$ 15.50 5.50 12.00 20.40 26.00 34.00 43.67 74.21 80.36 81.10
(c) $\checkmark$ $\checkmark$ $\checkmark$ 15.50 7.00 16.00 18.00 28.33 6.33 25.00 74.20 80.36 81.71
X-Boundary $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 17.50 7.50 16.00 10.40 16.67 5.33 15.00 74.17 80.52 81.10
4.3 Defense Performance on Reasoning Models
Recently, Jiang et al. (2025); Zhou et al. (2025) have highlighted significant safety risks in large reasoning models, such as DeepSeek-R1 (Guo et al., 2025). Notably, the reasoning outputs of R1 models pose even greater safety concerns than their final answers. In this subsection, we explore the implementation of X-Boundary to enhance the safety defenses of reasoning models. As shown in Table 3, X-Boundary effectively reduces the ASR of both single-turn and multi-turn attacks, while not causing severe over-refusal issues or a decline in reasoning ability in distilled reasoning models. Surprisingly, the CB appears to be less effective for reasoning models, possibly because of differences in their representation space compared to standard models. According to the theoretical insights in Section 3.3, X-Boundary reduces the difficulty of training and accelerates the training process. Therefore, X-Boundary has an advantage over CB. The effect of defense methods on the reasoning ability of R1 models is discussed in Appendix A.2.
4.4 Ablation Study
We conduct ablation studies on the impact of multi-turn defense data, boundary-safe data, and separate loss. The results are illustrated in Table 4. More results on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2 are shown in Appendix A.4.
Multi-turn defense data contribute to the reduction of ASR but intensify the over-refusal problem. With the multi-turn defense data described in Section 2.1 added into erase set, the ASR of ActorAttack is reduced from 63.00% to 15.50% on Qwen2.5-7B-Chat. However, the over-refusal rates in OR-Bench and PHTest increase about 30.00%. This highlights that over-refusal is a critical issue in multi-turn defense tasks, which cannot be overlooked and urgently requires resolution.
Boundary-safe data can partially mitigate the over-refusal issue. Boundary-safe QA pairs added to the retain set significantly reduce the over-refusal rate on OR-Bench and PHTest but show limited effectiveness on XSTest and OKTest. This may be because the boundary-safe QA pairs are synthesized by LLMs, leading to effectiveness on OR-Bench and PHTest, which also use synthetic data for testing. In contrast, the test queries in XSTest and OKTest are manually crafted and may differ in distribution from the synthetic data, making it difficult to achieve effective generalization.
Simply adjusting the size of boundary-safe data can not effectively balance ASR and over-refusal rate. Increasing the size of boundary-safe data can reduce the over-refusal rate, but it also leads to a sharp increase in ASR against multi-turn jailbreaks. Please see Appendix A.5 for more detailed results.
Separate loss can further reduce the over-refusal rate. Unlike simply adding boundary-safe data, separate loss markedly reduces the over-refusal rate on both manually crafted and synthetically constructed benchmarks. Since the boundary-safe data shares the same source as OR-Bench, simply adding data is sufficient to reduce the over-refusal rate to a very low level, leaving little room for separate loss to make a noticeable impact. However, in the other three benchmarks, separate loss further reduces the over-refusal rate by an average of 9.75% compared to adding boundary-safe data.
5 Related Work
Multi-turn attack.
Several studies have explored the safety risks in multi-turn dialogue scenarios (Wang et al., 2025; Tong et al., 2024). For instance, Li et al. (2024a) employs human red teamers to uncover vulnerabilities in LLMs when subjected to multi-turn attacks. Jiang et al. (2024) crafts 40 multi-turn scenarios in which malicious intent is concealed under the guise of preventing harm. Yu et al. (2024), Zhou et al. (2024b) and Liu et al. (2024b) generate multi-turn jailbreak queries by breaking down the original malicious query into multiple less harmful sub-questions. Ren et al. (2024b) captures multi-turn attack clues by modeling a network of semantically linked actors. Yang et al. (2024b) and Russinovich et al. (2024) dynamically adjust the attack query based on the contextual feedback from victim LLMs, gradually steering benign initial queries toward more harmful topics throughout the conversation. In this paper, we evaluate the defense robustness of existing methods and X-Boundary against three types of multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024), Crescendo (Russinovich et al., 2024).
Defenses for LLMs. Although defense methods for multi-turn jailbreak attacks are less explored in the literature, some existing approaches have proven effective against various single-turn attacks and have the potential to be adapted for multi-turn scenarios. These defense methods can be classified into the following categories: training LLMs to refuse harmful queries (Bai et al., 2022; Rafailov et al., 2024; Ouyang et al., 2022a; Yuan et al., 2024), training LLMs to prioritize safe instructions (Lu et al., 2024b; Wallace et al., 2024; Zhang et al., 2023), unlearning and editing harmful knowledge  (Lu et al., 2024a; Zhang et al., 2024b; Ren et al., 2024a; Qian et al., 2024a), prompt engineering (Xie et al., 2023; Zheng et al., 2024), and implementing input and output guardrails (Inan et al., 2023; Dubey et al., 2024) such as jailbreak detection (Hu et al., 2024a; Jain et al., 2023) input perturbation (Cao et al., 2023; Robey et al., 2023; Liu et al., 2024c). Several studies  (Li et al., 2024b; Zou et al., 2024, 2023; Qian et al., 2024b; Zhang et al., 2024a) also propose defense methods from the perspective of representation engineering, inspiring us to optimize LLMs in the representation space to strike a balance between defense robustness and LLM usability.
Decline in usability caused by defense methods. We assess the impact of defense methods on usability from two aspects: general capability degradation and over-refusal. General capability degradation, commonly known as the “alignment tax” (Ouyang et al., 2022b) phenomenon, has garnered widespread attention and has been extensively discussed in technical reports on LLMs (Dubey et al., 2024; Inan et al., 2023; Ren et al., 2024b; Li et al., 2024b; Hu et al., 2024b). Over-refusal refers to the unreasonable rejection of safe queries by LLMs (Varshney et al., 2023; Zhao et al., 2024; Zou et al., 2023; Arditi et al., 2024; Cao et al., 2024). Bianchi et al. (2023) discover that excessive safety-tuning makes LLMs refuse entirely safe prompts if they superficially resemble unsafe ones. Röttger et al. (2023), Shi et al. (2024), Cui et al. (2024), and An et al. (2024) employ linguistic techniques or automatic pipelines to generate seemingly unsafe prompts for evaluating LLMs’ over-refusal behavior.
6 Conclusion
In this paper, we comprehensively compare existing defense methods in multi-turn attack scenarios and reveal their shortcomings in balancing the robustness of defense and LLM usability. We analyze this issue from the perspective of LLMs’ feature space, and conclude that previous methods fail to learn a precise boundary that distinguishes safe and harmful representations without an explicit formulation. To address this issue, we propose the X-Boundary to push harmful representations away from safe representations through explicit loss functions and obtain a clear distinction boundary. Such distinction boundary enables the consequential removal of harmful representations without disrupting safe ones, thereby achieving a balance between robustness against multi-turn jailbreaks and LLM usability. We think that X-Boundary can offer more efficient and fine-grained defense for LLMs, complementing existing safety alignment techniques and ultimately improving the deployment of robust AI systems in real-world applications.
Impact Statement
This work aims to advance the field of large language models (LLMs) safety alignment by proposing X-Boundary, a method that maintains state-of-the-art performance in multi-turn jailbreak attack defenses while effectively mitigating the over-safety problem. We do not consider that this method will directly lead to severe negative consequences for societal development. However, we must be aware that malicious actors could exploit various approaches to induce LLMs to generate misleading or harmful content. Therefore, we expect that future research will focus on enhancing content moderation mechanisms and setting up ethical usage protocols to effectively reduce potential risks. From a constructive perspective, this method can significantly enhance the reliability, safety, and usability of LLMs.
Appendix A Additional Results
A.1 The compliance rates in HarmBench
We evaluate the basic safety level of the LLMs using HarmBench (Mazeika et al., 2024). Table 5 shows both existing defense methods and X-Boundary can effectively reduce the compliance rates to harmful requests in HarmBench without using attack methods.
Table 5: The compliance rates to harmful requests in HarmBench after using existing defense methods and X-Boundary.
Methods Llama-3- 8B-Instruct Qwen2.5- 7B-Chat Mistral-7B- Instruct-v0.2 Qwen2.5- 14B-Chat
Vanilla 11.67 26.25 56.67 15.83
SFT 1.25 5.42 7.08 1.25
DPO 0.83 2.92 13.33 5.83
GA 5.00 3.75 9.17 /
CB 1.67 4.58 13.33 4.58
X-Boundary 1.25 4.58 13.33 2.92
A.2 The effect of defense methods on the LLMs’ reasoning ability
Large reasoning models often rely on generating lengthy reasoning paths for inference. Therefore, we conducted a statistical analysis of the output length of large reasoning models employing various defense mechanisms. As shown in Table 6, while X-Boundary does not lead to a degradation in general capability, it results in shorter output lengths, which may indirectly impact reasoning performance. Exploring strategies to prevent the reduction in output length represents a promising direction for future research.
Table 6: Comparison of pass@1 accuracy and average output token length across different defense methods on reasoning model
Models Methods AIME2024 GPQA LiveCode
pass@1 Length (Avg.) pass@1 Length (Avg.) pass@1 Length (Avg.)
DeepSeek- R1-Distill- Llama-8B Vanilla 50.00 15672.07 50.00 8910.93 40.00 6457.43
SFT 44.95 13678.53 40.00 8699.93 35.10 6804.28
DPO 46.97 15716.27 50.00 8489.33 42.40 6301.96
CB 46.97 15488.23 46.97 9088.78 40.65 6479.9
X-Boundary 50.00 13310.90 50.00 8233.20 39.86 6498.04
DeepSeek- R1-Distill- Qwen-7B Vanilla 53.33 11046.63 48.99 8592.54 39.76 6683.22
SFT 46.67 13844.87 48.99 8176.29 36.44 6825.17
DPO 53.33 12063.57 50.00 8344.05 40.08 6694.74
CB 46.97 12609.93 46.97 8356.40 39.33 6536.76
X-Boundary 53.33 12959.73 50.51 8237.67 40.02 6583.29
A.3 The Trade-Off between Robustness and General Capability
Fig. 8 intuitively shows the trade-off between the ASR against multi-turn jailbreaks and the decline of general capability. As the training process advances, the ASR steadily decreases, while the decline in code and math capability progressively increases. X-Boundary lies in the lower-left corner of the plots, demonstrating that it achieves a win-win outcome with robust defense and strong general capability.
Figure 8: The trade-off between ASR of multi-turn jailbreak and general capability on Llama-3-8B-Instruct. The data points were collected by sampling and evaluating at every 100 training steps.
A.4 Ablation Study on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2
Through analyzing the results of ablation experiments in Table 7 and Table 8, we can obtain conclusions consistent with that in Section 4.4.
A.5 Effects of the Size of Boundary-Safe Data
Fig. 10 shows that as the boundary-safe data size increases, the over-refusal rate generally decreases, while ASR against multi-turn attacks tends to increase. Without the separate loss, when the boundary-safe data size reaches 500, the ASR hardly decreases, failing to achieve the purpose of enhancing multi-turn defense. This demonstrates that it is difficult to balance ASR and over-refusal rate simply by adjusting the boundary-safe data size.
A.6 Effects of Separate Loss and Boundary-Safe Data on the Representation Distribution
Fig. 11 shows that adding boundary-safe data to the retain set reduces the angle between boundary-safe representations after training and their original representations. Furthermore, under the effect of separate loss, this angle is further minimized. Meanwhile, the angle between boundary-safe representations and refusal representations increases, indicating that separate loss contribute to establish a clear distinction boundary.
Table 7: Ablation study on Llama-3-8B-Instruct. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
Models A B C D Multi-turn ASR (%) $\downarrow$ Over-refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 58.50 25.00 34.00 6.80 9.00 8.00 13.67 68.30 79.08 59.18
(a) $\checkmark$ 36.50 5.00 18.00 12.00 16.00 14.33 26.00 68.13 78.54 59.76
(b) $\checkmark$ $\checkmark$ 16.50 0.50 10.00 23.60 27.67 36.00 52.00 67.66 78.47 59.76
(c) $\checkmark$ $\checkmark$ $\checkmark$ 15.00 0.50 10.00 14.00 18.00 11.67 35.33 68.05 78.47 59.76
X-Boundary $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 16.50 1.00 10.00 8.40 14.00 8.00 28.66 67.94 78.47 59.76
Table 8: Ablation study on Mistral-7B-Instruct-v0.2. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
Models A B C D Multi-turn ASR (%) $\downarrow$ Over-refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 70.00 49.50 40.00 10.00 21.00 4.33 13.00 59.98 45.34 34.76
(a) $\checkmark$ 46.00 28.00 20.00 28.80 28.00 18.00 23.00 59.92 44.66 34.76
(b) $\checkmark$ $\checkmark$ 15.00 11.50 12.00 45.20 32.33 55.00 50.00 59.91 46.63 33.54
(c) $\checkmark$ $\checkmark$ $\checkmark$ 13.50 30.00 14.00 35.60 25.67 12.67 38.67 60.06 46.17 35.37
X-Boundary $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 16.00 13.50 14.00 19.20 23.33 10.33 26.33 59.83 45.34 36.59
Figure 10: The impact of boundary-safe data size on ASR and over-refusal rate without and with separate loss.
Figure 11: Visualization of effects of separate loss and boundary-safe data on the representation distribution. “Boundary-Safe” refers to the average representations of boundary-safe queries from OR-Bench along with their corresponding helpful responses. “refusal” refers to the average representations of boundary-safe queries from OR-Bench paired with refusal responses.
A.7 Details about Representation Visualization
To analyze safety-usability trade-off from the perspective of interpretability mechanism, we extract the feature representations from the 10th layer of Llama-3-8B-Instruct and visualize them using 2-dimensional t-SNE, as shown in Fig. 12.
Figure 12: Visualization of the representation distribution before and after implementing SFT, DPO, GA, and CB. “Harmful” and “boundary-safe” refer to the representations of harmful and boundary-safe queries along with their corresponding responses, respectively.
Appendix B Experimental Details
B.1 Construction of Multi-Turn Defense Dataset
We construct a multi-turn defense dataset based on SafeMTData. SafeMTData is derived from the circuit breaker training dataset, and carefully filtered to prevent data contamination with Harmbench. It includes harmful multi-turn queries generated by ActorAttack (Ren et al., 2024b), along with refusal responses to reject the harmful queries. To curate the harmful responses for DPO, GA, and CB, we use harmful multi-turn queries in SafeMTData to attack deepseek-chat (Liu et al., 2024a) and filter the harmful response using HarmBench classifier (Mazeika et al., 2024).
For SFT and DPO, following Ren et al. (2024b), we maintain a 1:2 ratio between the multi-turn defense data and instruction data, e.g., UltraChat (Ding et al., 2023). For CB, we add the filtered harmful responses and their corresponding single-turn queries into the CB dataset. The other data settings remain consistent with Zou et al. (2024). For GA, we follow (Zhang et al., 2024b) and use unlearning data, instruction data, and refusal data in a ratio of 5:5:1.
B.2 Training Details of Baselines
Multi-Turn SFT
For multi-turn SFT, we set the batch size to 1 with accumulation step 16. The training process was conducted for a total of 1 epoch. Optimization was performed using the AdamW optimizer, with the learning rate set to $5\times 10^{-4}$, ensuring stable and efficient model updates. The warm-up ratio and weight decay ratio are set to 0.05, 0.03. All training processes use Low-Rank Adaptation (LoRA) for parameter fine-tuning, where the rank $r$, scaling factor $\alpha$, and dropout rate are set to 16, 16, and 0.1, respectively. It takes about 40 minutes to train a Llama-3-8B-Instruct model on a single A100 80G GPU.
Multi-Turn DPO
For Multi-turn DPO, we use a learning rate of $1.0\times 10^{-5}$ with a cosine learning rate scheduler and a warm-up ratio of 0.1. We set the training epoch to 3 and the batch size to 1 with gradient accumulation steps of 8. All training processes use Low-Rank Adaptation (LoRA) for parameter fine-tuning with the rank $r$, scaling factor $\alpha$, and dropout rate set to 8, 16, and 0, respectively. We conducted all training processes on a single A100 80GB GPU.
Gradient Ascent
Following the experimental setting of Zhang et al. (2024b), we set the batch size to 11 with accumulation step 1, where the ratio of the three types of data in a batch is 5:5:1. We use the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and set the maximum epoch as 3. For Qwen2.5-7B-Chat and Llama-3-8B-Instruct, the coefficients of safe responses loss $\mathcal{L}_{s}$, general performance loss $\mathcal{L}_{g}$, and unlearning loss $\mathcal{L}_{h}$ are set to 0.5, 1.0, 0.3. For Mistral-7B-Instruct-v0.2, the loss coefficients are set to 0.25, 1.0, and 0.05, respectively. All training processes use Low-Rank Adaptation (LoRA) for parameter fine-tuning. For Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, we set the rank $r$, scaling factor $\alpha$, and dropout rate to 16, 16, 0.05. For Qwen2.5-7B-Chat, we conducted a grid search over the LoRA hyperparameters with $r\in\{8,16,32\}$ and $\alpha\in\{16,32,64\}$. We end up selecting $r=8$, $\alpha=64$, and a dropout rate of $0.05$. We linearly decay the learning rate and select the checkpoint after 1 epoch for evaluation. Training a Mistral-7B-Instruct-v0.2 model on a single A100 80GB GPU takes approximately 1 hour.
Circuit Breaker
We follow (Zou et al., 2024) to use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Chat and Qwen2.5-14B-Chat. We gather the feature representations from layers 10, 20, 30, and 40 to calculate circuit-breaking loss and inset LoRA adapter into all linear layers from 0 through 40. The loss coefficients are dynamically adjusted. The coefficients of circuit-breaking loss and retain loss are $c_{s}=\alpha(1-\frac{t}{\beta})$ and $c_{r}=\alpha\frac{t}{\beta}$, respectively. We set $\alpha$ as 5 on Mistral-7B-Instruct-v0.2 and 10 on other LLMs, $\beta$ as 300 on Mistral-7B-Instruct-v0.2 and Llama-3-8B-Instruct, 600 on Qwen2.5-7B-Chat, and 1200 on Qwen2.5-14B-Chat. Qwen2.5-14B-Chat is trained on for 360 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs is trained on for 180 steps with a batch size of 16 on 1 A100 GPU.
B.3 Training Details of X-Boundary
We use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Chat and Qwen2.5-14B-Chat. We set dynamic loss coefficients following (Zou et al., 2024), where $c_{r}=\alpha\frac{t}{\beta}$ and $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$. $\alpha,\beta$, and the target layers for calculating erase loss keep consistent with hyperparameters specified in Appendix B.2. We conduct a grid search on the size of boundary-safe data in a valid set in the range of [0,500], with a step of 50, selecting the size for Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.2, Qwen2.5-7B-Chat, and Qwen2.5-14B-Chat is 500, 200, 100, and 50, respectively. Qwen2.5-14B-Chat is trained for 260 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs are trained for 180 steps with a batch size of 16 on 1 A100 GPU.
B.4 Evaluations
Datasets.
We evaluate our approach on benchmarks covering multi-turn attacks, over-refusal, and general model capabilities:
Multi-Turn Attack
We employ three state-of-the-art multi-turn attack benchmarks. We adopt three state-of-the-art multi-turn attack benchmarks:
  • ActorAttack (Ren et al., 2024b): Emphasizes role-playing scenarios to gradually induce harmful behavior. The multi-turn queries in SafeMTData_Attack_600 (Ren et al., 2024b) are used to attack victim models, and HarmBench classifier (Mazeika et al., 2024) is used to judge whether the attack is successful.
  • RedQueen (Jiang et al., 2024): Focuses on dynamic prompt engineering with iterative refinements. We use the template of RedQueen to generate 600 test data based on HarmBench, and use HarmBench classifier as the judge model.
  • Crescendo (Russinovich et al., 2024): Includes gradually escalating attacks that push the model to produce harmful content over multiple turns. GPT-3.5-turbo is used as the attack model and GPT-4o is utilized as the judge model.
Over-Safety Assessment
We utilize four complementary datasets to measure over-refusal:
  • XSTest (Röttger et al., 2023): Examines model responses to boundary-case prompts involving sensitive but potentially valid information.
  • OKTest (Shi et al., 2024): Evaluates whether the model declines benign questions in real-world scenarios.
  • OR-Bench (Cui et al., 2024): Explicitly measures over-refusal rates on a suite of harmless queries.
  • PHTest: Comprises prompts that may look suspicious but are legitimately safe for the model to address.
General Capability
To ensure our method preserves the model’s general performance, we use:
  • MMLU: A broad measure of knowledge in diverse domains.
  • GSM8K: A math reasoning benchmark to test step-by-step problem solving.
  • HumanEval (Chen et al., 2021): Assesses code generation capability, crucial for real-world AI applications.
Evaluation Metrics.
To comprehensively assess our method, we adopt the following evaluation metrics:
  • Attack Success Rate (ASR): The proportion of attack attempts (single-turn or multi-turn) that successfully elicit harmful content from the model. Lower ASR indicates better robustness against jailbreaks.
  • Over-Refusal Rate : The fraction of benign prompts that the model incorrectly refuses to answer. A lower over-refusal rate signifies better usability.
  • General Performance : We measure the model’s utility on standard benchmarks (MMLU, GSM8K, HumanEval) to ensure that defensive measures do not degrade essential capabilities. A higher score indicates stronger performance on domain knowledge, reasoning, or code generation.
Appendix C Theoretical Analysis of X-Boundary
Proposition 1.
If $\phi_{\#}\mu$ is $(n,\Delta)$-clusterable, then for all $m\leq n(2\Delta)^{-2}$,
$\mathrm{Var}_{m}(\phi_{\#}\mu)<48\Delta.$ (1)
Given a distribution $\mu$, $(n,\Delta)$-clusterable means that $\textnormal{supp}(\mu)$ lies in the union of $n$ balls of radius at most $\Delta$.
Proof.
Proposition 1 in this paper is an application of Proposition 13 in (Weed & Bach, 2019).
Definition 1 ((Weed & Bach, 2019)).
A distribution $\mu$ is $(m,\Delta)$-clusterable if $\operatorname{supp}(\mu)$ lies in the union of $m$ balls of radius at most $\Delta$.
Proposition 2 (Proven in (Weed & Bach, 2019)).
If $\mu$ is $(n,\Delta)$-clusterable, then for all $m\leq n(2\Delta)^{-2p}$,
$\mathbb{E}_{S\sim\mu^{m}}[{\mathcal{W}}_{p}^{p}(\mu,\mu_{S})]\leq(9^{p}+3)\sqrt{\frac{n}{m}}.$ (2)
According to the triangle inequality, we have
$\mathrm{Var}_{m}(\phi_{\#}\mu)=\mathbb{E}_{S,\tilde{S}\sim\mu^{m}}[{\mathcal{W}}_{1}(\phi_{\#}\mu_{S},\phi_{\#}\mu_{\tilde{S}})]\leq 2\mathbb{E}_{S\sim\mu^{m}}[{\mathcal{W}}_{p}^{p}(\mu,\mu_{S})]\leq 24\sqrt{\frac{n}{m}}.$ (3)
In this way, for all $m\leq n(2\Delta)^{-2p}$, we have
$\mathrm{Var}_{m}(\phi_{\#}\mu)\leq 24\sqrt{\frac{n}{m}}<48\Delta.$ (4)
Appendix D Case Study
In this section, we showcase a range of examples to demonstrate the practical utility of the X-Boundary in mitigating over-safety and its robustness against multi-turn attacks.
Cases of Over-Refusal In Fig. 13 and Fig. 14, we present two examples from OKTest and OR-Bench, comparing the responses of Llama-3-8B-Instruct with different safety strategies. X-Boundary provides accurate and useful responses to boundary-safe queries, e.g., “Is there an effective program in place to educate children about the dangers of molestation? ”, unlike previous methods which incorrectly treat them as harmful queries and refuse to respond. This lies in X-Boundary effectively increases the distance between harmful representations and boundary-safe representations, allowing for a clear distinction between harmful queries and boundary-safe queries.
Cases of Multi-Turn Attack In Fig. 15, Fig. 16, and Fig. 17, we present three examples taken from ActorAttack, RedQueen, and Crescendo, respectively. X-Boundary effectively defends against various types of multi-turn jailbreaks, demonstrating its robustness.
Cases of Safety Defense on Reasoning Model In Fig. 18, with X-Boundary, dangerous content generated in the reasoning process is automatically converted into gibberish. Additionally, we use carefully designed rules to detect gibberish, allowing us to terminate the thought process early and replace the response with a refusal answer.
Figure 13: Over-Safety example from Llama-3-8B-Instruct using various defense methods on OR-Bench.
Figure 14: Over-Safety example from Llama-3-8B-Instruct using various defense methods on OKTest.
Figure 15: Multi-Turn Attack example from Llama-3-8B-Instruct on Actorattack.
Figure 16: Multi-Turn Attack example from Llama-3-8B-Instruct on RedQueen.
Figure 17: Multi-Turn Attack example from Qwen2.5-7B-Chat on Crescendo.
Figure 18: Safety defense example on reasoning models.